Two sound waves travel out from a common point and have frequencies 40Hz and 30Hz. What is the phase difference between them after 0.8 seconds? 

Grade:11

1 Answers

Ankur Sarkar
5 years ago
​In one second the frequency difference is(40-30)=10
​In 0.8 s,the frequency difference(the difference in the number of vibration made) is 10*0.8=8
​for one cycle the phase difference is 2*3.14
​for 8 vibrations the phase difference is 2*3.14*8=16*3.14